# Retinex-MEF: Retinex-based Glare Effects Aware Unsupervised Multi-Exposure Image Fusion

### v1

@@ 3.4 Inference
Benefiting from the separation of exposure and reflectance components, our objective is to ensure that our fusion results remain independent of specific exposure levels, thereby allowing for flexible adjustments to global illumination. To achieve this goal, we have developed a controllable exposure fusion function, which is described as follows:
@@
In this function, $x\in[0,1]$ denotes the pre-adjustment exposure level, while $k\in[0,1]$ indicates the desired increase in exposure, serving as an adjustable parameter. Consequently, a higher $k$ value leads to a brighter image. Behavior of $f(x)$ across varying $k$ values is depicted in Fig. 6 (a).
This function possesses several features conducive to adjusting illumination: 1) It is monotonically increasing with respect to both $x$ and $k$, enabling a range of adjustments for varying $k$ values and simplifying to the identity mapping $f(x)=x$ when $k=0.5$. Additionally, the function fulfills the condition $f(x;k)+f(1-x;1-k)=1$, as depicted in Fig. 6 (b). This indicates that when the adjustment parameters are $k$ and $1-k$, respectively, the change in pixel value from $x$ to $f(x)$ is directly counterbalanced by the change from $1-x$ to $f(1-x)$, aligning with the symmetrical nature of exposure adjustments. To facilitate more nuanced exposure adjustments, we propose a reformulation of Eq.(14) as follows:
@@
In our fusion experiments and visualizations, we fix $k$ at 0.5, indicating that the illumination adjustment is governed by Eq.(14). For the exposure adjustment and homogeneous extreme exposure fusion experiments, we strictly follow the fusion criteria outlined in Eq.(16). During these adjustments, we define the exposure level $E$ as the average exposure across the entire image, which is calculated as follows:
@@ 4.1 Setup
In our experiments, the network is trained for 80 epochs using the Adam optimizer with a batch size of 8. The hyperparameters ${\alpha_{1},\alpha_{2},\alpha_{3},\alpha_{4}}$ are set to ${0.5,0.1,0.2,0.1}$ to ensure the consistency of loss magnitudes and appropriate prioritization. The initial learning rate is established at $10^{-4}$ and halved every 20 epochs. All experiments are performed on a PC equipped with a single NVIDIA RTX 3090 GPU.
Our method was compared with advanced fusion techniques such as CUNet [5], SDNet [54], DeFusion [24], CRMEF [29], DDMEF [43], MEFLUT [13], and FILM [60]. These fusion techniques were evaluated using metrics such as human visual perception ($Q_{cb}$), normalized mutual information (NMI), nonlinear correlation information entropy ($Q_{ncie}$), the Structural Similarity Index Measure (SSIM) for multi-exposure image fusion, Peak Signal-to-Noise Ratio (PSNR), and the correlation coefficient (CC). The performance was validated using two publicly accessible multi-exposure image fusion datasets SICE [3] and MEFB [56]. A total of 300 image sequences are extracted from the SICE dataset for training and 100 pairs for testing. Furthermore, we incorporated an additional validation set comprising 30 image pairs from the MEFB dataset to further substantiate the model’s performance under varied conditions.
@@ 4.2 Underexposure-Overexposure Fusion
Fig. 5 illustrates our method’s visual comparison on two datasets, displaying superior detail preservation in both overexposed and underexposed areas. In the first indoor scene, details on wall paintings and color layers are richer and clearer than other methods, with a clear outside window scene. In the outdoor scene, textures of roads and buildings are well-preserved. Details like distant trees and railings are notably clearer, with other methods showing blurring. Our glare modeling contributes to a balanced color distribution, preventing over-saturation and color biases, and enhancing visual comfort. More results are displayed in the supplementary material.
@@ 4.6 Ablation Study
Our unsupervised method’s performance is significantly influenced by the choice of loss functions, as demonstrated in the ablation studies shown in Fig. 10 and Tab. 2. In Experiment I, applying the reconstruction loss directly to $\hat{R}$ instead of the glare-affected ${R}$ resulted in color degradation. The omission of the smoothness loss in Experiment II led to a reflection component that captured color but lost structural and textural details. Without the initialization loss in Experiment III, the reflection component inadequately preserved colors, darkening the fused image. In Experiment IV, lacking suppression loss caused $\hat{R}$ to align with the glare-influenced $R$, affecting it with overexposure glare. Finally, Experiment V showed that without the consistency loss, $\hat{R}$ functioned merely as an intermediate feature, failing to produce a fused image. The declines in performance with each modification, as quantitative results affirm the effectiveness of our method’s loss functions. More visualization results are displayed in the supplementary material.
@@ 5 Conclusion
To counteract the effects of extreme exposure in source images on multi-exposure image fusion performance, we introduce an unsupervised framework based on Retinex theory. This framework effectively separates and independently refines the common scene reflectance components from the unique exposure levels of each image. By simulating the glare effects in overexposed regions, we safeguard against degradation in these areas, thereby preserving the integrity of common reflectance component extraction through the use of bidirectional constraints. Furthermore, we have developed a controllable exposure fusion function that adapts to various image exposure levels, enabling our model to skillfully blend and adjust exposure level, with flexibility extending beyond producing a single fused image. Experiments on various datasets, including exposure adjustment and fusion of images with abnormal exposures, demonstrate both the effectiveness and flexibility of our model.

### v2

@@ 3.4 Inference
The separation of exposure and reflectance components provides a principled foundation for manipulating scene illumination without altering its intrinsic content. Building upon prior works that leverage adjustment curves for brightness and contrast modulation [12, 65, 55, 14], we aim to decouple the fusion result from specific exposure levels, thereby enabling flexible illumination control. To achieve this, we introduce a controllable exposure fusion function:
@@
Here, $x\in[0,1]$ denotes the input exposure, and the parameter $k\in[0,1]$ directly controls the mapping intensity. The behavior of this function across varying $k$ values is depicted in Fig. 6 (a).
[…]
@@
In our fusion experiments and visualizations, we fix $k$ at 0.5, indicating that the illumination adjustment is governed by Eq. (14). For exposure adjustment and homogeneous extreme exposure fusion experiments, we strictly follow the fusion criteria outlined in Eq. (16). During these adjustments, we define the exposure level $E$ as the average exposure across the entire image, which is calculated as follows:
@@ 4.1 Setup
In our experiments, the network is trained for 80 epochs using the Adam optimizer with a batch size of 8. The hyperparameters ${\alpha_{1},\alpha_{2},\alpha_{3},\alpha_{4}}$ are set to 0.5,0.1,0.2,0.1{0.5,0.1,0.2,0.1}0.5 , 0.1 , 0.2 , 0.1 to ensure the consistency of loss magnitudes and appropriate prioritization. The initial learning rate is established at $10^{-4}$ and halved every 20 epochs. All experiments are performed on a PC equipped with a single NVIDIA RTX 3090 GPU.
Our method is compared with advanced fusion techniques such as CUNet [9], SDNet [67], DeFusion [33], CRMEF [40], DDMEF [53], MEFLUT [18], and FILM [76]. These fusion techniques are evaluated using metrics such as human visual perception ($Q_{cb}$), normalized mutual information (NMI), nonlinear correlation information entropy ($Q_{ncie}$), the Structural Similarity Index Measure (SSIM) for MEF, Peak Signal-to-Noise Ratio (PSNR), and the correlation coefficient (CC). The performance is validated using two publicly accessible multi-exposure image fusion datasets SICE [5] and MEFB [69]. A total of 300 image sequences are extracted from the SICE dataset for training and 100 pairs for testing. Furthermore, we incorporate an additional validation set comprising 30 image pairs from the MEFB dataset to further substantiate the model’s performance under varied conditions.
@@ 4.2 Underexposure-Overexposure Fusion
Fig. 5 illustrates our method’s visual comparison on two datasets, displaying superior detail preservation in both overexposed and underexposed areas. In the first indoor scene, details on wall paintings and color layers are richer and clearer than other methods, with a clear outside window scene. In the outdoor scene, the textures of roads and buildings are well-preserved. Details like distant trees and railings are notably clearer, with other methods showing blurring. Our glare modeling contributes to a balanced color distribution, preventing over-saturation and color biases, and enhancing visual comfort. More results are displayed in the supplementary material.
@@ 4.6 Ablation Study
The performance of our unsupervised method is critically dependent on the choice of loss functions, as demonstrated by the ablation studies in Fig. 10 and Tab. 2. In Exp. I, applying the $\mathcal{L}_{recon}$ directly to $\hat{R}$ instead of the glare-affected ${R}$ reduces the model to a traditional Retinex decomposition, causing a whitish color cast in the fused results. In Exp. II, omitting $\mathcal{L}_{smooth}$ causes the reflection component to capture color but lose structural and textural details. Without the initialization loss (Exp. III), the reflection component fails to preserve color information adequately, which darkens the fused image. In Exp. IV, the absence of $\mathcal{L}_{suppress}$ causes $\hat{R}$ to align with the glare-influenced $R$, leaving glare effects unresolved. Finally, Exp. V shows that without $\mathcal{L}_{consist}$, $\hat{R}$ functions merely as an intermediate feature and fails to produce a meaningful fused image. These performance declines across all variants, supported by quantitative results, affirm the necessity and effectiveness of each proposed loss function. Additional visual results are provided in the supplementary material.
@@ 5 Conclusion
To counteract the effects of extreme exposure in source images on multi-exposure image fusion performance, we introduce an unsupervised framework based on Retinex theory. This framework effectively separates and independently refines the common scene reflectance components from the unique exposure levels of each image. By simulating the glare effects in overexposed regions, we safeguard against degradation in these areas, thereby preserving the integrity of common reflectance component extraction through the use of bidirectional constraints. Furthermore, we design a parameterized illumination fusion strategy that adapts to various image exposure levels, enabling our model to skillfully blend and adjust exposure level, with flexibility extending beyond producing a single fused image. Experiments on various datasets, including exposure adjustment and abnormal exposure pairs fusion validate the superior effectiveness and unique flexibility of our model.
Acknowledgement
This work has been supported by the National Natural Science Foundation of China under Grant 12201497 and 12371512, Young Talent Fund of Xi’an Association for Science and Technology under Grant 0959202513207.